# From No-Regret to Strategically Robust Learning in Repeated Auctions

### v1

@@ 2.2 Repeated auction game with no-regret learners
    The auctioneer chooses an auction format $\mathcal{M}^{(t)}=(\mathbf{x}^{(t)},\mathbf{p}^{(t)})$ by specifying an allocation function $\mathbf{x}^{(t)}:B^{n}\to\Delta([n])$ and a payment function $\mathbf{p}^{(t)}:B^{n}\to[0,1]^{n}$.
@@
We let $\textnormal{Reg}_{i}$ denote bidder $i$’s regret––the difference between the cumulative utility of the best fixed strategy in hindsight and bidder $i$’s cumulative utility, i.e.,
@@ 2.3 Discussion on the model assumptions
Assumption on auction formats.
Our setting is in line with previous works by Braverman et al. (2018) and Kumar et al. (2024). The main distinction is the class of auction formats available to the auctioneer. Specifically, Braverman et al. (2018) impose no restrictions on auction formats. Kumar et al. (2024) focus exclusively on first-price auctions with reserve prices. Our assumption on auction formats (Assumption 2.4) strikes a middle ground that captures most auctions deployed in practice, including first-price, second-price, and all-pay auctions with reserve prices, posted-price auctions, and Myerson’s optimal auction. ∎
@@
In contrast to the standard no-regret learning literature, where randomization over strategies is typically necessary to guarantee no regret, our repeated auction setting does not require bidders to employ randomization when choosing their bidding strategies. Moreover, the auctioneer selects the auction format after observing bidders’ bidding strategies, which seems to make no-regret learning impossible. Nonetheless, we note that there is intrinsic randomness in bidders’ private values, which are sampled after the auction format is chosen. This randomness can be used by an algorithm to achieve no regret and strategic robustness. ∎
@@ 3 The quantile strategy space
In this section, we introduce the quantile strategy space, which will be used by our meta-algorithm. We note that our quantile strategy space is a variant of the one [^7] To be precise, this refers to the decision space of the concave formulation in Kumar et al. (2024, Section 3). used in Kumar et al. (2024), and more generally, quantile-based representations have been widely studied in the auction literature (see e.g., Hartline (2013, Chapter 3) and the references therein).
@@ 3.2 Quantile strategies and quantile utility functions
Now we introduce quantile strategies and reformulate each bidder’s no-regret learning problem on the quantile strategy space. Formally, our quantile strategy space is simply a probability simplex.
@@
$s_{i,\bm{\pi}}(v)=\begin{cases}0&\textrm{if }v\in[0,F_{i}^{-1}(\pi_{1})]\\ \frac{j-1}{K}&\textrm{if }v\in(F_{i}^{-1}(\sum_{\ell=1}^{j-1}\pi_{\ell}),F_{i}^{-1}(\sum_{\ell=1}^{j}\pi_{\ell})]\textrm{ for }j\in\{2,\dots,K+1\},\end{cases}$ (6)
@@
If a bidder $i\in[n]$ chooses a quantile strategy $\bm{\pi}$ at round $t\in[T]$ of the repeated auction game and bids according to the monotone bidding strategy $s_{i,\bm{\pi}}$, then the resulting utility is
@@
where $u_{i}^{(t)}$ is defined in (3). We call $q_{i}^{(t)}$ the quantile utility function. In Lemma 3.5, we reformulate bidder $i$’s regret on the quantile strategy space, using quantile utility functions.
@@ 3.3 Concavity of quantile utility functions
For any $i\in[n]$ and $t\in[T]$, the quantile utility function $q_{i}^{(t)}:\Delta([K+1])\to[-1,1]$, as defined in Eq. (7), is concave. Moreover, for all $k\in[K+1]$, the partial derivative of $q_{i}^{(t)}(\bm{\pi})$ with respect to $\pi_{k}$ can be expressed as
$\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}=\Big(\sum_{j=k}^{K}\big(x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)\big)\cdot F_{i}^{-1}\big(\sum_{\ell=1}^{j}\pi_{\ell}\big)\Big)+x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big),$ (9)
and we have that $\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}\in[-1,1]$ under Assumption 2.4-i.
Proof.
We first expand the function $q_{i}^{(t)}(\bm{\pi})$ as follows,
$q_{i}^{(t)}(\bm{\pi})$ $=u^{(t)}_{i}(s_{i,\bm{\pi}})$ (By (7))
$=\sum_{j\in[K+1]}\int_{F_{i}^{-1}(\sum_{\ell=1}^{j-1}\pi_{\ell})}^{F_{i}^{-1}(\sum_{\ell=1}^{j}\pi_{\ell})}x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)\cdot v_{i}-p^{(t)}_{i}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)dF_{i}(v_{i})$ (By Eq. (6))
$=\sum_{j\in[K+1]}\int_{\sum_{\ell=1}^{j-1}\pi_{\ell}}^{\sum_{\ell=1}^{j}\pi_{\ell}}x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)\cdot F_{i}^{-1}(z)-p^{(t)}_{i}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)dz.$ (10)
Then, we define the partial-sum functions $\tau_{j}(\bm{\pi}):=\sum_{\ell=1}^{j}\pi_{\ell}$ for all $j\in[K+1]$, and we define an intermediate function $\sigma_{i}^{(t)}:[0,1]^{K+1}\to[-1,1]$ as follows [^8] We note that $\sigma_{i}^{(t)}$ is a generalization of bidders’ utility functions in Kumar et al. (2024, Eq. (3)) from first-price auctions to arbitrary auctions.,
$\sigma_{i}^{(t)}(y_{1},\dots,y_{K+1}):=\sum_{j\in[K+1]}\int_{y_{j-1}}^{y_{j}}x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)\cdot F_{i}^{-1}(z)-p^{(t)}_{i}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)dz,$
where we let $y_{0}:=0$. We notice that $q_{i}^{(t)}(\bm{\pi})=\sigma_{i}^{(t)}(\tau_{1}(\bm{\pi}),\dots,\tau_{K+1}(\bm{\pi}))$ by Eq. (3.3), and moreover, for each $j\in[K+1]$, $\tau_{j}$ is a linear function of $\bm{\pi}$.
Concavity of $q_{i}^{(t)}$
Since linear transformations preserve concavity (Boyd and Vandenberghe, 2004, Section 3.2), to show that $q_{i}^{(t)}$ is concave, it suffices to prove that $\sigma_{i}^{(t)}$ is concave. To this end, we observe that the partial derivatives of $\sigma_{i}^{(t)}$ with respect to $y_{j}$, for all $j\in[K+1]$, are given by
$\frac{\partial\sigma_{i}^{(t)}}{\partial y_{j}}$ $=\big(x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)\big)\cdot F_{i}^{-1}(y_{j})-\big(p^{(t)}_{i}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)\big),\,\,\forall\,j\in[K],$
$\frac{\partial\sigma_{i}^{(t)}}{\partial y_{K+1}}$ $=x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(1,\mathbf{b}_{-i}^{(t)}\big).$ (11)
We notice that for all $j\in[K]$, the partial derivative $\frac{\partial\sigma_{i}^{(t)}}{\partial y_{j}}$ depends only on $y_{j}$, and $\frac{\partial\sigma_{i}^{(t)}}{\partial y_{K+1}}$ is constant. Thus, to prove that $\sigma_{i}^{(t)}$ is concave, it suffices to show that for all $j\in[K]$, $\frac{\partial\sigma_{i}^{(t)}}{\partial y_{j}}$ is a non-increasing function of $y_{j}$. We observe that $F_{i}^{-1}(y_{j})$ is non-decreasing in $y_{j}$ since $F_{i}$ is a CDF, and that $x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)$ is non-positive because $x_{i}^{(t)}$ is non-decreasing by Assumption 2.4-i. It follows that $\frac{\partial\sigma_{i}^{(t)}}{\partial y_{j}}$ is a non-increasing function of $y_{j}$ for all $j\in[K]$, which implies that $\sigma_{i}^{(t)}$, and hence $q_{i}^{(t)}$, is concave.
Gradient of $q_{i}^{(t)}$
Next, for each $k\in[K+1]$, we compute the partial derivative of $q_{i}^{(t)}$ with respect to $\pi_{k}$:
$\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}$ $=\sum_{j=1}^{K+1}\frac{\partial\sigma_{i}^{(t)}}{\partial\tau_{j}}\cdot\frac{\partial\tau_{j}}{\partial\pi_{k}}=\sum_{j=k}^{K+1}\frac{\partial\sigma_{i}^{(t)}}{\partial\tau_{j}}$
$\quad\quad+x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(1,\mathbf{b}_{-i}^{(t)}\big)$
(By Eq. (3.3))
$=\bigg(\sum_{j=k}^{K}\big(x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)\big)\cdot F_{i}^{-1}(\tau_{j}(\bm{\pi}))\bigg)+x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)$
$\qquad\text{(By a telescoping sum)},$
which establishes Eq. (9). To bound $\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}$, we notice that each term $x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)$ is non-positive (by Assumption 2.4-i) and $F_{i}^{-1}(\tau_{j}(\bm{\pi}))\in[0,1]$. Hence, we have that
Combining this with Eq. (9), we obtain that
$\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}$ $\leq x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)\leq 1,$
$\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}$ $\geq x_{i}^{(t)}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)+x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)$
$=x_{i}^{(t)}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)\geq-1,$
which completes the proof of the lemma. ∎
@@ 4 From no-regret to strategically robust learning
Input : No-regret learning algorithm $\mathcal{A}$, and the CDF $F_{i}$ of bidder $i$’s value distribution
@@ 3   Algorithm $\mathcal{A}$ outputs a distribution $\bm{\pi}^{(i,t)}\in\Delta([K+1])$;
6   Observe the utility function $u_{i}^{(t)}$ (as defined in Eq. (3)) and the corresponding quantile utility function $q_{i}^{(t)}$ (as defined in Eq. (7));
@@
By Lemma 3.6, the reward vectors $\nabla q_{i}^{(t)}(\bm{\pi}^{(i,t)})$ provided to the input algorithm $\mathcal{A}$ are in the range $1,1]^{K+1}$, which is consistent with the standard online learning setting (Definition 4.1). Hence, by our assumption in the lemma statement, algorithm $\mathcal{A}$ guarantees that
@@ 4.1 Auxiliary auctions
For each round $t\in[T]$, given the auction format $\mathcal{M}^{(t)}=(\mathbf{x}^{(t)},\mathbf{p}^{(t)})$ selected by the auctioneer and bidders’ monotone bidding strategies $s_{i}^{(t)}$ chosen by Algorithm 1, the auxiliary auction $\tilde{\mathcal{M}}^{(t)}:=(\tilde{\mathbf{x}}^{(t)},\tilde{\mathbf{p}}^{(t)})$ is specified by an allocation function $\tilde{\mathbf{x}}^{(t)}:[0,1]^{n}\to\Delta([n])$ and a payment function $\tilde{\mathbf{p}}^{(t)}:[0,1]^{n}\to[0,1]^{n}$. Specifically, the allocation function $\tilde{\mathbf{x}}^{(t)}$ is defined as
@@
Intuitively, the auxiliary auction $\tilde{\mathcal{M}}^{(t)}$ can be viewed as a variant of $\mathcal{M}^{(t)}$ whose allocation function $\mathbf{x}^{(t)}$ is “stretched” by the monotone bidding strategies $s_{i}^{(t)}$ (see Figure 2 for an illustration). Moreover, we note that $\tilde{\mathcal{M}}^{(t)}$ has a continuous bid space $[0,1]$, which is consistent with Myerson’s classic setting introduced in Section 2.1. In Claim 4.5, we show that $\tilde{\mathcal{M}}^{(t)}$ is IC and IR.
@@
Figure 2: This figure illustrates an example of an auxiliary auction $\tilde{\mathcal{M}}$ in the single-bidder case. On the left is the allocation function $x$ for the bidder in the original auction $\mathcal{M}$. On the right is the allocation function $\tilde{x}$ for the bidder in the auxiliary auction $\tilde{\mathcal{M}}$, where $\tilde{x}(v)=x(s(v))$ is obtained by composing $x$ with a monotone bidding strategy $s(v):=\frac{1}{2}\cdot\mathds{1}(v\geq\frac{1}{2})$.
@@
In Claim 4.6, we derive a useful expression for the expected revenue $\textnormal{Rev}_{\tilde{\mathcal{M}}^{(t)}}(\mathcal{D})$ obtained by the auxiliary auction $\tilde{\mathcal{M}}^{(t)}$ in the classic auction setting.
@@
We clarify that in Algorithm 1, the bid vector $\mathbf{b}_{-i}^{(t)}$ is determined by the bidders’ quantile/bidding strategies and their random private values $\mathbf{v}_{-i}^{(t)}$ at round $t$. Hence, the expectation in the above equation is taken with respect to the randomness of $\mathbf{v}_{-i}^{(t)}$.

### v2

@@ 2.2 Repeated auction game with no-regret learners
    The auctioneer chooses an auction format $\mathcal{M}^{(t)}=(\mathbf{x}^{(t)},\mathbf{p}^{(t)})$ by specifying an allocation function $\mathbf{x}^{(t)}:B^{n}\to\mathcal{X}([n])$ and a payment function $\mathbf{p}^{(t)}:B^{n}\to[0,1]^{n}$.
@@
We let $\textnormal{Reg}_{i}$ denote bidder $i$’s regret—the difference between the cumulative utility of the best fixed strategy in hindsight and bidder $i$’s cumulative utility, i.e.,
@@ 2.3 Discussion on the model assumptions
Assumption on auction formats and bidders’ feedback.
Our setting is in line with previous works by Braverman et al. (2018) and Kumar et al. (2024). The main distinction is the class of auction formats available to the auctioneer. Specifically, Braverman et al. (2018) impose no restrictions on auction formats. Kumar et al. (2024) focus exclusively on first-price auctions with reserve prices. Our assumption on auction formats (Assumption 2.4) strikes a middle ground that captures most auctions deployed in practice, including first-price, second-price, and all-pay auctions with reserve prices, posted-price auctions, and Myerson’s optimal auction.
Another difference concerns bidders’ feedback. In our general setup, we consider the full-information setting, rather than the more general bandit-feedback setting considered in Braverman et al. (2018). In our meta-algorithm, it suffices that, at each round $t\in[T]$, each bidder $i\in[n]$ can observe $x_{i}^{(t)}\big(b,\mathbf{b}_{-i}^{(t)}\big)$ and $p_{i}^{(t)}\big(b,\mathbf{b}_{-i}^{(t)}\big)$ for all $b\in B$. For example, in a first-price auction, a bidder can infer these quantities given the highest competing bid, which coincides with the feedback setting studied in Kumar et al. (2024). ∎
@@
In contrast to the standard no-regret learning literature, where randomization over strategies is typically necessary to guarantee no regret, our repeated auction setting does not require bidders to employ randomization when choosing their bidding strategies. Moreover, the auctioneer selects the auction format after observing bidders’ bidding strategies, which seems to make no-regret learning impossible. Nonetheless, we note that there is intrinsic randomness in bidders’ private values, which are sampled after the auction format is chosen. This randomness can be used by an algorithm to achieve no regret and strategic robustness, as demonstrated by Kumar et al. (2024). ∎
@@ 3 The quantile strategy space
In this section, we introduce the quantile strategy space, which will be used by our meta-algorithm. We note that our quantile strategy space is a variant of the one [^7] To be precise, this refers to the decision space of the concave formulation in Kumar et al. (2024, Section 3), and a related formulation appears in Kumar and Mouchtaki (2025). used in Kumar et al. (2024), and more generally, quantile-based representations have been widely studied in the auction literature (see e.g., Hartline (2013, Chapter 3) and the references therein).
@@ 3.2 Quantile strategies and quantile utility functions
Now we introduce quantile strategies and reformulate each bidder’s online optimization problem on the quantile strategy space. Formally, our quantile strategy space is simply a probability simplex.
@@
$s_{i,\bm{\pi}}(v)=\begin{cases}0&\textrm{if }v\in[0,F_{i}^{-1}(\pi_{1})]\\ \frac{j-1}{K}&\textrm{if }v\in(F_{i}^{-1}(\sum_{\ell=1}^{j-1}\pi_{\ell}),F_{i}^{-1}(\sum_{\ell=1}^{j}\pi_{\ell})]\textrm{ for }j\in\{2,\dots,K\}\\ 1&\textrm{if }v\in(F_{i}^{-1}(\sum_{\ell=1}^{K}\pi_{\ell}),1]\end{cases},$ (6)
@@
If a bidder $i\in[n]$ chooses a quantile strategy $\bm{\pi}$ at round $t\in[T]$ of the repeated auction game and bids according to the monotone bidding strategy $s_{i,\bm{\pi}}$, then the resulting utility is [^8] We note that $q_{i}^{(t)}$ is independent of $\pi_{K+1}$, since $s_{i,\bm{\pi}}$ does not depend on $\pi_{K+1}$ by Eq. (6).
@@
where $u_{i}^{(t)}$ is defined in Eq. (3). We call $q_{i}^{(t)}$ the quantile utility function. In Lemma 3.5, we reformulate bidder $i$’s regret on the quantile strategy space, using quantile utility functions.
@@ 3.3 Concavity of quantile utility functions
For any $i\in[n]$ and $t\in[T]$, the quantile utility function $q_{i}^{(t)}:\Delta([K+1])\to[-1,1]$, as defined in Eq. (7), is concave. Moreover, we have that $\frac{\partial q_{i}^{(t)}}{\partial\pi_{K+1}}=0$, and for all $k\in[K]$, the partial derivative of $q_{i}^{(t)}(\bm{\pi})$ with respect to $\pi_{k}$ can be expressed as
$\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}=\Big(\sum_{j=k}^{K}\big(x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)\big)\cdot F_{i}^{-1}\big(\sum_{\ell=1}^{j}\pi_{\ell}\big)\Big)+p_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big),\,\forall\,k\in[K].$ (9)
Furthermore, for all $k\in[K+1]$, we have that $\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}\in[-2,1]$ under Assumption 2.4-i.
The proof is provided in Section A.3.
@@ 4 From no-regret to strategically robust learning
Input : Online learning algorithm $\mathcal{A}$, and the CDF $F_{i}$ of bidder $i$’s value distribution $\mathcal{D}_{i}$
@@ 3   Algorithm $\mathcal{A}$ outputs a distribution $\bm{\pi}^{(i,t)}\in\Delta([K+1])$;
6   Observe $x_{i}^{(t)}\big(b,\mathbf{b}_{-i}^{(t)}\big)$ and $p_{i}^{(t)}\big(b,\mathbf{b}_{-i}^{(t)}\big)$ for all $b\in B$, which (together with $F_{i}$) are sufficient to compute the gradient $\nabla q_{i}^{(t)}(\bm{\pi}^{(i,t)})$ (see Lemma 3.6);
@@
By Lemma 3.6, the reward vectors $\nabla q_{i}^{(t)}(\bm{\pi}^{(i,t)})$ provided to the input algorithm $\mathcal{A}$ are in the range $2,1]^{K+1}$, which is consistent with the standard online learning setting (Definition 4.1). Hence, by our assumption in the lemma statement, algorithm $\mathcal{A}$ guarantees that
@@ 4.1 Auxiliary auctions
For each round $t\in[T]$, given the auction format $\mathcal{M}^{(t)}=(\mathbf{x}^{(t)},\mathbf{p}^{(t)})$ selected by the auctioneer and bidders’ monotone bidding strategies $s_{i}^{(t)}$ chosen by Algorithm 1, the auxiliary auction $\tilde{\mathcal{M}}^{(t)}:=(\tilde{\mathbf{x}}^{(t)},\tilde{\mathbf{p}}^{(t)})$ is specified by an allocation function $\tilde{\mathbf{x}}^{(t)}:[0,1]^{n}\to\mathcal{X}([n])$ and a payment function $\tilde{\mathbf{p}}^{(t)}:[0,1]^{n}\to[0,1]^{n}$. Specifically, the allocation function $\tilde{\mathbf{x}}^{(t)}$ is defined as
@@
Intuitively, the auxiliary auction $\tilde{\mathcal{M}}^{(t)}$ can be viewed as a variant of $\mathcal{M}^{(t)}$ whose allocation function $\mathbf{x}^{(t)}$ is “stretched” by the monotone bidding strategies $s_{i}^{(t)}$ (see Figure 2 for an illustration).
@@
Figure 2: This figure illustrates an example of an auxiliary auction $\tilde{\mathcal{M}}$ in the single-bidder case. On the left is the allocation function $x$ for the bidder in the original auction $\mathcal{M}$. On the right is the allocation function $\tilde{x}$ for the bidder in an auxiliary auction $\tilde{\mathcal{M}}$, where $\tilde{x}(v)=x(s(v))$ is obtained by composing $x$ with a monotone bidding strategy $s(v):=\frac{1}{2}\cdot\mathds{1}(v\geq\frac{1}{2})$.
Moreover, we note that $\tilde{\mathcal{M}}^{(t)}$ has a continuous bid space $[0,1]$, which is consistent with Myerson’s setting introduced in Section 2.1. In Claim 4.5, we show that $\tilde{\mathcal{M}}^{(t)}$ is IC and IR.
@@
In Claim 4.6, we derive a useful expression for $\textnormal{Rev}_{\tilde{\mathcal{M}}^{(t)}}(\mathcal{D})=\operatorname*{\mathbb{E}}_{\mathbf{v}\sim\mathcal{D}}\big[\sum_{i\in[n]}\tilde{p}_{i}^{(t)}(\mathbf{v})\big]$, namely, the expected revenue obtained by the auxiliary auction $\tilde{\mathcal{M}}^{(t)}$ in the classic auction setting.
@@
We clarify that in Algorithm 1, the bid vector $\mathbf{b}_{-i}^{(t)}$ is determined by the bidders’ quantile/bidding strategies and their private values $\mathbf{v}_{-i}^{(t)}$ at round $t$. Hence, the expectation on the R.H.S. of the above equation is taken over the randomness of $\mathbf{v}_{-i}^{(t)}$, which induces randomness in $\mathbf{b}_{-i}^{(t)}$.